OpenStudy (rock_mit182):

The diameters of the main rotor and tail rotor of a single-engine helicopter are 7.60 m and 1.02 m, respectively. The respective rotational speeds are 450 rev/min and 4 138 rev/min. Calculate the speeds of the tips of both rotors. Compare these speeds with the speed of sound, 343 m/s.

OpenStudy (sidsiddhartha):

for the first one 450 rev/min=47.1rad/sec=omega and radius=7.60/2=3.80 speed of first rotor=v=r*(omega)=3.80*47.1=179 m/s to comapre with speed of sound divide 179 by 343=52.2%

OpenStudy (rock_mit182):

thanks